# Nonparametric Estimation of Matching Efficiency and Elasticity on a Private On-the-Job Search Platform: Evidence from Japan, 2014-2024

### v4

@@ Abstract
I analyze proprietary data from BizReach (2014–2024) to estimate the matching function for high-skill workers on a private on-the-job search platform using Lange and Papageorgiou (2020) nonparametric approach. Comparing it to Hello Work, I find that matching efficiency on the private platform is both more volatile and higher, reflecting its growing popularity. Matching elasticity with respect to users is around 0.75, while for vacancies, it reaches 1.0, suggesting a more balanced elasticity than Hello Work. The study also uncovers industry-level heterogeneity, highlighting differences in matching dynamics across sectors.
@@ 1 Introduction
On-the-job search plays a crucial role in labor reallocation, driving wage and productivity improvements (Moscarini and Postel-Vinay 2017). In the U.S., job-to-job transitions account for one-third to one-half of all hires (Faberman et al. 2022). While Japan has historically been characterized by long-term employment stability, recent trends indicate a growing prevalence of on-the-job search. The proportion of employed individuals changing jobs remains relatively low at 4.8% in 2023, but the share of workers actively seeking new opportunities reached a record high of 15.3%, marking the tenth consecutive year of growth. [^1] According to the Labor Force Statistics Office of the Statistics Bureau, Ministry of Internal Affairs and Communications. Despite the increasing significance of on-the-job search in labor markets worldwide, empirical research on its matching efficiency and elasticity—especially in the context of private job platforms—remains scarce. This stands in contrast to the more extensive literature on job search behavior among unemployed individuals.
This paper fills this gap by examining long-term trends in on-the-job search platforms for high-skill workers across multiple labor markets. Using proprietary aggregate data from BizReach, a leading private online job scouting platform, I estimate the matching function and recover matching efficiency and elasticity through a nonparametric approach developed by Lange and Papageorgiou (2020). Unlike conventional job boards where workers actively apply for vacancies, BizReach allows registered job seekers to receive scouting messages from companies and headhunters searching for specialized talent. As of July 2024, over 2.58 million employed and self-employed workers were registered on the platform, underscoring its relevance in analyzing the dynamics of on-the-job search.
@@ 1.1 Related literature
First, it adds to the empirical literature on the estimation of the matching function, a foundational component in macroeconomic models. Using a novel nonparametric approach developed by Lange and Papageorgiou (2020), I examine trends in matching efficiency in Japanese labor markets via an online job scouting platform. This method enables the identification and estimation of the matching function without imposing the standard independence assumption between matching efficiency and search efforts from either side of the labor market. This approach accommodates multiple types of job seekers. Lange and Papageorgiou (2020) highlight the positive correlation between efficiency and market structure variables like labor market tightness, which introduces a positive bias in vacancy elasticity estimates unless unobserved matching efficacy is accounted for. In traditional Cobb-Douglas matching function models, this unobserved factor is often ignored, resulting in potentially biased elasticity estimates. [^2] For example, Petrongolo and Pissarides (2001) summarize early aggregate studies using the Cobb-Douglas matching function, finding the match elasticity with respect to unemployment to be in the range of 0.5 to 0.7. In the context of Japan, Otani (2024) update the earlier findings of Kano and Ohta (2005), Kambayashi and Ueno (2006), and others, comparing results with international findings like Bernstein et al. (2022).
@@
Second, this paper relates to the Japanese labor market literature, particularly in the context of the public off-the-job search platform, Hello Work. [^3] See the survey of Miyamoto (2025) for an overview of labor markets in Japan. Otani (2024) estimate matching efficiency and mismatch in Japan’s Hello Work platform between 1972 and 2024, showing a declining trend in matching efficiency consistent with decreasing job and worker finding rates. The match elasticity with respect to unemployment is found to be between 0.5 and 0.9, while the elasticity concerning vacancies ranges from 0.1 to 0.4. In comparison, Kanayama and Otani (2024) apply a similar nonparametric method to estimate matching efficiency and elasticity in a privately-operated spot-worker platform, contrasting these findings with Hello Work’s part-time data. My paper complements and extends these findings by focusing on a high-skill, employed worker platform, filling an important gap in the literature.
The Japanese government’s efforts to enhance labor market flexibility have significantly contributed to the increase in on-the-job search activities. Initiatives such as the 2018 Work Style Reform Laws, aimed at reducing overwork, improving work-life balance, and promoting diverse career trajectories, have encouraged workers to explore new career opportunities without the societal stigma previously associated with job changes. These reforms have reshaped attitudes toward job mobility and created a more flexible labor market environment, enabling workers to transition between roles more easily while still employed (Yamamoto 2019, Owan 2017). This policy context underscores the relevance of quantitative studies on private on-the-job search platforms, such as those examined in this paper.
@@
Second, this paper contributes to the literature on on-the-job search, a key aspect of labor search theory since the 1970s (Burdett 1978). Recent theoretical models, including those by Cahuc et al. (2006), Eeckhout and Lindenlaub (2019), and Bagger and Lentz (2019), emphasize the role of search effort in on-the-job search and its connection to job ladder dynamics. Empirical research, such as Mueller (2010) and Ahn and Shao (2017), has relied on data from the American Time Use Survey (ATUS) to document on-the-job search behaviors. However, due to the limitations of ATUS in capturing search outcomes, there is a gap in evaluating the efficiency of on-the-job search. Notably, Faberman et al. (2022) and Roussille and Scuderi (2023) provide crucial insights, with Faberman et al. (2022) focusing on the relationship between search effort and outcomes and Roussille and Scuderi (2023) exploring wage markdown using data from Hired.com. However, these studies lack long-term macro-level insights into matching function, efficiency, and elasticity. The proprietary data in this paper offers a unique advantage by allowing for the evaluation of matching efficiency in the on-the-job search labor market via a private platform.
@@
Third, this paper contributes to the expanding literature on online job search platforms. The analysis of job matching within real-world market institutions has gained prominence due to the increasing availability of data from online job platforms (Autor 2019). [^4] Examples include studies like Kuhn and Skuterud (2004), Kuhn and Mansour (2014), and Kroft and Pope (2014), which focus on worker status, while others, such as Kuhn and Shen (2013), Hershbein and Kahn (2018), Brown and Matsa (2016), and Azar et al. (2020), focus solely on vacancy data. Moreover, research like Banfi and Villena-Roldan (2019), Marinescu and Rathelot (2018), Marinescu and Wolthoff (2020), and Azar et al. (2022) incorporates both worker and vacancy information. Much of the literature emphasizes application-level or vacancy-level behavior to assess search behavior and wage elasticity. For instance, Faberman and Kudlyak (2019) leverage proprietary application-level data from an online job search engine to explore the relationship between search intensity and duration, primarily focusing on lower-skill, hourly jobs for employed and unemployed workers. Similarly, Kambayashi et al. (2025) estimate elasticities of application, interview attendance, and offer acceptance relative to posted wages using detailed process-level data from private job search and matching intermediary platforms in Japan, which became the most significant recruitment channel in 2023, as shown in Figure 1. In contrast, this paper adopts a broader macro-level perspective, evaluating the overall efficiency of the private matching platform. To the best of my knowledge, this is the first paper to estimate matching efficiency and elasticity in an online job scouting platform, offering relatively long-term insights into private online job search trends, complementing the micro-level studies.
@@ 2.1 Data source
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for month-level aggregate data from January 2014 to April 2024 to examine trends in matching unemployed workers with vacancies via Japan’s public employment platform, Hello Work. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which regularly publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work plays a crucial role in Japan’s labor market by providing government-operated employment counseling, job placement services, and vocational training. It has been extensively used for estimating traditional Cobb-Douglas matching functions, as seen in studies like Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018), as well as nonparametric estimation Otani (2024). In this study, I focus on full-time workers to ensure consistency for comparison across different datasets. The chosen period provides a consistent timeframe for comparison with the following platform data.
@@ 3 Model
Our primary focus lies in analyzing matching efficiency and matching elasticity with respect to the number of registered workers and vacancies in the labor market, as facilitated by an online job scouting platform operated by a private firm in Japan. A matching function derived from search models plays a pivotal role in labor economics. [^5] See Pissarides (2000), Petrongolo and Pissarides (2001), and Rogerson et al. (2005) for further reference. The matching function operates on the premise of random search from both sides of the labor market, where job seekers represent labor supply and recruiters represent labor demand. Conceptually, this paper examines two independent matching functions on the off-the-job and on-the-job search labor markets. Examining interdependence between two matching functions theoretically is out of the scope of this paper.
@@
To estimate the matching function and recover matching efficiency, I adopt the novel approach proposed by Lange and Papageorgiou (2020). The paper highlights two critical issues: the endogeneity of matching efficiency (Borowczyk-Martins et al. 2013) and the overly restrictive nature of the Cobb-Douglas specification, which assumes fixed matching elasticity. To address these limitations, Lange and Papageorgiou (2020) propose a nonparametric identification and estimation framework for matching efficiency under specific conditions that will be discussed later.
@@
I identify the matching function and the unobserved, time-varying matching efficiency, $A$. First, I assume that $V$ and $A$ are conditionally independent given $U$, i.e., $A\perp V\mid U$ [^6] Matching models often incorporate a free entry condition, whereby firms continue to post job vacancies as long as the expected return is positive. This condition implies that the number of vacancies is endogenously determined and may depend on current matching efficiency, potentially violating the exogeneity assumption (Borowczyk-Martins et al. 2013). While Lange and Papageorgiou (2020) additionally incorporate search effort data (Mukoyama et al. 2018) and a recruitment intensity index data (Davis et al. 2013) to attempt to address this endogeneity issue using time-use survey data—which is typically unavailable in most contexts—Otani (2024) instead demonstrates through Monte Carlo simulations that the resulting bias is economically negligible, with estimated parameters deviating by less than 5% from the time-varying true values, and benefit of capturing time-varying efficiency is substantial. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ exhibits constant returns to scale (CRS). [^7] To align with the original model of Matzkin (2003), the function $H=m(AU,V)$ can be reformulated as $H/U=m(A,V/U)$ under CRS, where $H/U$ and $V/U$ are the job-finding rate and market tightness, respectively. These two assumptions are commonly used in the literature. Applying the nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) demonstrates that the joint distribution $G(H,U,V)$ identifies $F(A,U)$ and the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point, denoted as $A_{0}$, within the support of $(A,U,V)$. [^8] In Otani (2024), I report finite sample performance and extend the methodology through Monte Carlo simulation. Simulation results with a sample size of $T=50$ indicate that the sample size in this paper is sufficient for accurately recovering matching efficiency. For practical issues and application in different contexts, see Brancaccio et al. (2020b), which applies this approach to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023), and Otani (2025) which applies this approach to marriage markets.
@@ 4 Estimation
Following Lange and Papageorgiou (2020), I begin by estimating $F(A_{0}|U)$ across the support of $U$. To achieve this, I use the distribution of hires conditional on users, $U$, and observed vacancies, $V$. Specifically, we have:
@@ 5.1 Matching efficiency and elasticity in the platform
Figure 3 presents a comparison of matching efficiency and matching elasticity between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. I set the reference period to $t=0$ corresponding to January 2014, and normalize matching efficiency by setting $A_{0}=100$. In the Hello Work platform, as depicted in panel (a), matching efficiency remains relatively stable around the baseline (normalized to January 2014) until 2021 but sharply declines thereafter. Panel (b) illustrates the matching elasticities with respect to unemployment and vacancies. The elasticity with respect to unemployment remains consistently lower, around 0.4, while the elasticity with respect to vacancies shows a gradual increase, reaching values near 1.0 by 2022. This suggests that changes in vacancies have a larger impact on the number of matches than changes in unemployment, and the assumption of constant returns to scale has been seemingly violated after 2020.101010The estimated elasticities differ from those reported in Otani (2024), largely due to the different time horizons considered. Specifically, Otani (2024) includes data from 1972 to 2024, a period that encompasses various economic booms and busts, which likely captures a broader range of labor market dynamics and affects the elasticity estimates.
@@ 6 Conclusion
Additionally, industry-level heterogeneity is evident across both platforms, reflecting differing labor market dynamics by sector. However, while this paper provides key insights into the matching function for on-the-job searches, the analysis may not fully represent the broader on-the-job search labor market, particularly for non-high-skill workers in Japan. Moreover, the standard assumption of homogeneity among workers and vacancies may overlook important nuances. Future research should focus on expanding this analysis to other private platforms and exploring individual-level behavior, as discussed in studies like Kambayashi et al. (2025) and Roussille and Scuderi (2023), to provide a more comprehensive understanding of labor market dynamics.

### v5

@@ Abstract
I analyze proprietary data from BizReach (2014–2024) to estimate the matching function for high-skill workers on a private on-the-job search platform using lange2020beyond nonparametric approach. Comparing it to Hello Work, I find that matching efficiency on the private platform is both more volatile and higher, reflecting its growing popularity. Matching elasticity with respect to users is around 0.75, while for vacancies, it reaches 1.0, suggesting a more balanced elasticity than Hello Work. The study also uncovers industry-level heterogeneity, highlighting differences in matching dynamics across sectors.
@@ 1 Introduction
On-the-job search plays a crucial role in labor reallocation, driving wage and productivity improvements (moscarini2017relative). In the U.S., job-to-job transitions account for one-third to one-half of all hires (faberman2022job). While Japan has historically been characterized by long-term employment stability, recent trends indicate a growing prevalence of on-the-job search. The proportion of employed individuals changing jobs remains relatively low at 4.8% in 2023, but the share of workers actively seeking new opportunities reached a record high of 15.3%, marking the tenth consecutive year of growth. [^1] According to the Labor Force Statistics Office of the Statistics Bureau, Ministry of Internal Affairs and Communications. Despite the increasing significance of on-the-job search in labor markets worldwide, empirical research on its matching efficiency and elasticity—especially in the context of private job platforms—remains scarce. This stands in contrast to the more extensive literature on job search behavior among unemployed individuals.
This paper fills this gap by examining long-term trends in on-the-job search platforms for high-skill workers across multiple labor markets. Using proprietary aggregate data from BizReach, a leading private online job scouting platform, I estimate the matching function and recover matching efficiency and elasticity through a nonparametric approach developed by lange2020beyond. Unlike conventional job boards where workers actively apply for vacancies, BizReach allows registered job seekers to receive scouting messages from companies and headhunters searching for specialized talent. As of July 2024, over 2.58 million employed and self-employed workers were registered on the platform, underscoring its relevance in analyzing the dynamics of on-the-job search.
@@ 1.1 Related literature
First, it adds to the empirical literature on the estimation of the matching function, a foundational component in macroeconomic models. Using a novel nonparametric approach developed by lange2020beyond, I examine trends in matching efficiency in Japanese labor markets via an online job scouting platform. This method enables the identification and estimation of the matching function without imposing the standard independence assumption between matching efficiency and search efforts from either side of the labor market. This approach accommodates multiple types of job seekers. lange2020beyond highlight the positive correlation between efficiency and market structure variables like labor market tightness, which introduces a positive bias in vacancy elasticity estimates unless unobserved matching efficacy is accounted for. In traditional Cobb-Douglas matching function models, this unobserved factor is often ignored, resulting in potentially biased elasticity estimates. [^2] For example, petrongolo2001looking summarize early aggregate studies using the Cobb-Douglas matching function, finding the match elasticity with respect to unemployment to be in the range of 0.5 to 0.7. In the context of Japan, otani2024nonparametric update the earlier findings of kano2005estimating, kambayashi2006vacancy, and others, comparing results with international findings like bernstein2022matching.
@@
This paper also relates to the Japanese labor market literature, particularly in the context of the public off-the-job search platform, Hello Work. [^3] See the survey of miyamoto2025macroeconomic for an overview of labor markets in Japan. otani2024nonparametric estimate matching efficiency and mismatch in Japan’s Hello Work platform between 1972 and 2024, showing a declining trend in matching efficiency consistent with decreasing job and worker finding rates. The match elasticity with respect to unemployment is found to be between 0.5 and 0.9, while the elasticity concerning vacancies ranges from 0.1 to 0.4. In comparison, kanayama2024nonparametric apply a similar nonparametric method to estimate matching efficiency and elasticity in a privately-operated spot-worker platform, contrasting these findings with Hello Work’s part-time data. My paper complements and extends these findings by focusing on a high-skill, employed worker platform, filling an important gap in the literature.
The Japanese government’s efforts to enhance labor market flexibility have significantly contributed to the increase in on-the-job search activities. A prominent example is the 2018 Work Style Reform Acts, which aimed to reduce overwork, improve work-life balance, and promote diverse career trajectories. While these reforms were not directly designed to foster job transitions, measures such as the promotion of side jobs and secondary employment have encouraged workers to explore new career pathways, helping to reduce the societal stigma previously associated with job changes.
At the same time, the expansion of educational training benefits through the Employment Insurance System has provided workers with greater opportunities to reskill and acquire advanced competencies. Together with structural pressures from persistent labor shortages since the 2010s, these developments have further reshaped attitudes toward mobility and created an environment more conducive to job transitions. The growth of the job transition market should therefore be understood not as the product of the Work Style Reform Acts alone, but rather as the outcome of a broader set of institutional reforms and structural changes. This policy context underscores the relevance of quantitative studies on private on-the-job search platforms, such as those examined in this paper.
@@
Second, this paper contributes to the literature on on-the-job search, a key aspect of labor search theory since the 1970s (burdett1978theory). Recent theoretical models, including those by cahuc2006wage, eeckhout2019unemployment, and bagger2019empirical, emphasize the role of search effort in on-the-job search and its connection to job ladder dynamics. Empirical research, such as mueller2010job and ahn2017precautionary, has relied on data from the American Time Use Survey (ATUS) to document on-the-job search behaviors. However, due to the limitations of ATUS in capturing search outcomes, there is a gap in evaluating the efficiency of on-the-job search. Notably, faberman2022job and roussille2023bidding provide crucial insights, with faberman2022job focusing on the relationship between search effort and outcomes and roussille2023bidding exploring wage markdown using data from Hired.com. However, these studies lack long-term macro-level insights into matching function, efficiency, and elasticity. The proprietary data in this paper offers a unique advantage by allowing for the evaluation of matching efficiency in the on-the-job search labor market via a private platform.
@@
Third, this paper contributes to the expanding literature on online job search platforms. The analysis of job matching within real-world market institutions has gained prominence due to the increasing availability of data from online job platforms (autor2019studies). [^4] Examples include studies like kuhn2004internet, kuhn2014internet, and kroft2014does, which focus on worker status, while others, such as kuhn2013gender, hershbein2018recessions, brown2016boarding, and azar2020concentration, focus solely on vacancy data. Moreover, research like banfi2019high, marinescu2018mismatch, marinescu2020opening, and azar2022estimating incorporates both worker and vacancy information. Much of the literature emphasizes application-level or vacancy-level behavior to assess search behavior and wage elasticity. For instance, faberman2019intensity leverage proprietary application-level data from an online job search engine to explore the relationship between search intensity and duration, primarily focusing on lower-skill, hourly jobs for employed and unemployed workers. Similarly, kambayashi2025decomposing estimate elasticities of application, interview attendance, and offer acceptance relative to posted wages using detailed process-level data from private job search and matching intermediary platforms in Japan, which became the most significant recruitment channel in 2023, as shown in Figure 1. In contrast, this paper adopts a broader macro-level perspective, evaluating the overall efficiency of the private matching platform. To the best of my knowledge, this is the first paper to estimate matching efficiency and elasticity in an online job scouting platform, offering relatively long-term insights into private online job search trends, complementing the micro-level studies.
@@ 2.1 Data source
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for month-level aggregate data from January 2014 to April 2024 to examine trends in matching unemployed workers with vacancies via Japan’s public employment platform, Hello Work. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which regularly publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work plays a crucial role in Japan’s labor market by providing government-operated employment counseling, job placement services, and vocational training. It has been extensively used for estimating traditional Cobb-Douglas matching functions, as seen in studies like kano2005estimating, kambayashi2006vacancy, sasaki2008matching, and higashi2018spatial, as well as nonparametric estimation otani2024nonparametric. In this study, I focus on full-time workers to ensure consistency for comparison across different datasets. The chosen period provides a consistent timeframe for comparison with the following platform data.
@@ 3 Model
Our primary focus lies in analyzing matching efficiency and matching elasticity with respect to the number of registered workers and vacancies in the labor market, as facilitated by an online job scouting platform operated by a private firm in Japan. A matching function derived from search models plays a pivotal role in labor economics. [^5] See pissarides2000equilibrium, petrongolo2001looking, and rogerson2005search for further reference. The matching function operates on the premise of random search from both sides of the labor market, where job seekers represent labor supply and recruiters represent labor demand. Conceptually, this paper examines two independent matching functions on the off-the-job and on-the-job search labor markets. Examining interdependence between two matching functions theoretically is out of the scope of this paper.
@@
To estimate the matching function and recover matching efficiency, I adopt the novel approach proposed by lange2020beyond. The paper highlights two critical issues: the endogeneity of matching efficiency (borowczyk2013accounting) and the overly restrictive nature of the Cobb-Douglas specification, which assumes fixed matching elasticity. To address these limitations, lange2020beyond propose a nonparametric identification and estimation framework for matching efficiency under specific conditions that will be discussed later.
@@
I identify the matching function and the unobserved, time-varying matching efficiency, $A$. First, I assume that $V$ and $A$ are conditionally independent given $U$, i.e., $A\perp V\mid U$. [^6] Matching models often incorporate a free entry condition, whereby firms continue to post job vacancies as long as the expected return is positive. This condition implies that the number of vacancies is endogenously determined and may depend on current matching efficiency, potentially violating the exogeneity assumption (borowczyk2013accounting). While lange2020beyond additionally incorporate search effort data (mukoyama2018job) and a recruitment intensity index data (davis2013establishment) to attempt to address this endogeneity issue using time-use survey data—which is typically unavailable in most contexts—otani2024nonparametric instead demonstrates through Monte Carlo simulations that the resulting bias is economically negligible, with estimated parameters deviating by less than 5% from the time-varying true values, and benefit of capturing time-varying efficiency is substantial. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ exhibits constant returns to scale (CRS). [^7] To align with the original model of matzkin2003nonparametric, the function $H=m(AU,V)$ can be reformulated as $H/U=m(A,V/U)$ under CRS, where $H/U$ and $V/U$ are the job-finding rate and market tightness, respectively. These two assumptions are commonly used in the literature. Applying the nonparametric identification results of matzkin2003nonparametric, Proposition 1 of lange2020beyond demonstrates that the joint distribution $G(H,U,V)$ identifies $F(A,U)$ and the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point, denoted as $A_{0}$, within the support of (A,U,V)(A,U,V). [^8] In otani2024nonparametric, I report finite sample performance and extend the methodology through Monte Carlo simulation. Simulation results with a sample size of $T=50$ indicate that the sample size in this paper is sufficient for accurately recovering matching efficiency. For practical issues and application in different contexts, see brancaccio2020guide, which applies this approach to estimate a matching function in a trade model (brancaccio2020geography, brancaccio2023search), and otani2025nonparametric which applies this approach to marriage markets.
@@ 4 Estimation
Following lange2020beyond, I begin by estimating $F(A_{0}|U)$ across the support of $U$. To achieve this, I use the distribution of hires conditional on users, $U$, and observed vacancies, $V$. Specifically, we have:
@@ 5.1 Matching efficiency and elasticity in the platform
Figure 3 presents a comparison of matching efficiency and matching elasticity between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. I set the reference period to $t=0$ corresponding to January 2014, and normalize matching efficiency by setting $A_{0}=100$. In the Hello Work platform, as depicted in panel (a), matching efficiency remains relatively stable around the baseline (normalized to January 2014) until 2021 but sharply declines thereafter. Panel (b) illustrates the matching elasticities with respect to unemployment and vacancies. The elasticity with respect to unemployment remains consistently lower, around 0.4, while the elasticity with respect to vacancies shows a gradual increase, reaching values near 1.0 by 2022. This suggests that changes in vacancies have a larger impact on the number of matches than changes in unemployment, and the assumption of constant returns to scale has been seemingly violated after 2020.101010The estimated elasticities differ from those reported in otani2024nonparametric, largely due to the different time horizons considered. Specifically, otani2024nonparametric includes data from 1972 to 2024, a period that encompasses various economic booms and busts, which likely captures a broader range of labor market dynamics and affects the elasticity estimates. Also, see the Cobb-Douglas specification results in Appendix A.
@@ 6 Conclusion
Additionally, industry-level heterogeneity is evident across both platforms, reflecting differing labor market dynamics by sector. However, while this paper provides key insights into the matching function for on-the-job searches, the analysis may not fully represent the broader on-the-job search labor market, particularly for non-high-skill workers in Japan. Moreover, the standard assumption of homogeneity among workers and vacancies may overlook important nuances. Future research should focus on expanding this analysis to other private platforms and exploring individual-level behavior, as discussed in studies like kambayashi2025decomposing and roussille2023bidding, to provide a more comprehensive understanding of labor market dynamics.
Appendix A Appendix
Table 1 presents country–month level estimates using Hello Work data to illustrate model comparisons. In Panel (a), the country–month level Cobb–Douglas estimates appear implausible: the constant term suggests negative matching efficiency, and the elasticities sum to well above unity, violating the constant-returns-to-scale condition. These issues likely stem from aggregation at the national level, which masks regional and occupational heterogeneity and inflates the responsiveness of hires to unemployment and vacancies. In addition, the small number of observations reduce stability. By contrast, Panel (b) shows that estimation at finer levels of aggregation—such as prefecture–month or occupation–month—yields more plausible results with fixed effects, with reasonable elasticities summing closer to one. Taken together, these patterns suggest that the implausible results should not be interpreted as evidence against the constant returns to scale and positive efficiency per se, but rather highlight that the estimates are sensitive to the level of aggregation and the sample period used.
While the results for BizReach are confidential and cannot be reported, the estimates are more reasonable than the Hello Work results. Moreover, otani2024nonparametric demonstrates through Monte Carlo simulations that the nonparametric approach yields lower bias and Root Mean Squared Error (RMSE) and is robust to functional form misspecification, further supporting the usefulness of this framework.
Table 1 compares model performance across datasets. In the case of BizReach, where matching efficiency is time-varying, the nonparametric approach provides superior performance in terms of RMSE and Mean Absolute Error (MAE). For Hello Work, however, where matching efficiency appears relatively stable over 2014–2024, the constant-efficiency OLS specification delivers a better fit by suppressing fluctuations. Nevertheless, such stability is itself revealed only when allowing for nonparametric estimation, underscoring the continued value of the nonparametric approach.
Table 1: Estimation Results Using Cobb-Douglas Matching Function: Hello Work Data
Dependent var log(H)
Constant -56.673
log(U) 2.688
log(V) 2.123
Num.Obs. 124
R2 0.661
R2 Adj. 0.656
RMSE 0.16
(1) (2)
Dependent var log(H) log(H)
log(U) 0.266 0.510
(0.019) (0.013)
log(V) 0.319 0.546
(0.014) (0.010)
FE pref, month occupation, month
Num.Obs. 6392 8908
R2 0.987 0.992
R2 Adj. 0.987 0.992
Note: FE: fixed effects. In Panel (a), I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for country-month-level aggregate data from January 2014 to April 2024 to examine trends in matching unemployed full-time workers with vacancies via Japan’s public employment platform, Hello Work. In Panel (b), I extract results from otani2024nonparametric using the prefecture-month-level and occupation-month-level data from January 2012 to March 2023.
Table 2: Model Performance Comparison
Cobb Douglas Nonparametric
RMSE 0.1834 0.1821
MAE 0.1387 0.1237
Cobb Douglas Nonparametric
RMSE 0.1639 0.2006
MAE 0.1359 0.1579
Note: RMSE = Root Mean Squared Error, MAE = Mean Absolute Error.